Your firm is expected to generate free cash flows (FCFs) of $160 million annually in perpetuity. The first FCF will occur one year from today. The corporate tax rate is 20%. The comparable firm for you company, called A-Co. has a total market value of $600 million. A-Co. has expected FCFs of $36M per year, forever. A-Co. is an all-equity company. a. Assume your company will never take on any debt. What is the value of your company? b. Now assume that you will keep a constant interest coverage ratio. Because of this, your company will have $640M in debt today (time 0). The cost of debt for your company is 3%. What is the value of your company if you keep a constant interest coverage ratio?
Answer #1

(a) Your Firm FCF = $ 160 million, Co A: FCF = $ 36 million and Value = $ 600 million

Therefore, Your Firm Value = (600/36) x 160 = $ 2666.67 million

(b) When the firm decides to maintain a constant interest coverage ratio, it implies that the ratio of the operating income to the interest expense is constant. This in turn would mean that the debt value remains constant if the interest rate remains fixed at 3%.

Hence, Value generated by constant tax shield = Debt Level x Tax Rate = 640 x 0.2 = $ 128 million

Firm Value = 2666.67 + 128 = $ 2794.67 million